OpenStudy (anonymous):

Find the inverse of the following function. Is the inverse a function? y=(x-8)^2 I have an idea on how to do it I'm just not really sure and I'd like to know steps on how to do it.

OpenStudy (anonymous):

Switch y and x in the equation and solve for y.

OpenStudy (anonymous):

Step 1. x= (y-8) ^2 I have switched x and y

OpenStudy (anonymous):

Step 2..x= (y-8)(y-8). Can you figure (y-8)(y-8) ?

OpenStudy (anonymous):

So it would be y^2-16y+64? Then what? Solve for y?

OpenStudy (anonymous):

I would switch the x and the y... then solve it so, x=(y-8)^2 take the root of both sides now, +-root(x) = y - 8 add 8 to boths and y = +=-rootx - 8 this webside it .. nevermind

OpenStudy (anonymous):

My bad... Ok we have the inverse as x=(y-8)^2. Now we have to isolate the y. So take the sqaure root of both sides gives you what?

OpenStudy (anonymous):

The square root of both sides give you +-rootx=y-8? I'm not sure.

OpenStudy (anonymous):

\[\sqrt{x}=y-8 You mean this?

OpenStudy (anonymous):

\[\sqrt{x}=y-8\]

OpenStudy (anonymous):

Yeah pretty much.

OpenStudy (anonymous):

Now isolate y on one side of the equation and you have your inverse function. To solve any inverse just swap x and y in the original equation and solve for y.

OpenStudy (anonymous):

Yeah, that's right. I had that on my homework and it was that.

OpenStudy (anonymous):

Did you come up with an answer?

OpenStudy (anonymous):

So in the end it would be y=+-√x-8?

OpenStudy (anonymous):

\[y=\sqrt{x}+8\]

OpenStudy (anonymous):

Add eight to both sides

OpenStudy (anonymous):

Ohhh alright. Thanks. But I wouldn't be +-√x?

OpenStudy (anonymous):

good observation @Josh422, if you take the square root of an expression you get a plus and a minus evaluation.

OpenStudy (anonymous):

Oh wow thank you!

OpenStudy (anonymous):

so what does that tell you? Is it a function?

OpenStudy (anonymous):

Well wouldn't I have to graph it first to tell if it's one to one?

OpenStudy (anonymous):

You're kind of right, a function and the graph of a function are basically the same thing. However, you can choose a more analytic approach. Maybe you have already had circles? and their definition? Circles clearly are no functions. Why? Because we give up on right uniqueness and therefore give up on functionality.

OpenStudy (anonymous):

I'm fairly sure the answer is \[y=\sqrt{x} +8\]

OpenStudy (anonymous):

Not sure where the +/- comes from.

OpenStudy (anonymous):

Well yeah if it's the root then it should be plus or minus. Let's say "x" was 4. The square root of 4 can also be -2, as Negative two times Negative two would be 4.

OpenStudy (anonymous):

You cannot have a negative number under a radical. You are thinking more in line with x^2 = 4..Where x can be 2 or -2.

OpenStudy (anonymous):

@nelsonjedi, truth to be told, the question above lacks of description. It makes as good as no sense to talk about inverse functions when there is not even a domain and a codomain defined. Remember that you can only invert a function if it is bijective, meaning injective and surjective. Hence the above function is neither at the moment. If we suppose it is bijective, then your answer is right and we can drop the plus or minus sign. But since this statement was not made, we cannot do such a thing and we would be better off by keeping the plus or minus sign.

OpenStudy (anonymous):

Well thank you both for the help, I don't know how to give you badges to help you for helping me. But I'll have too go soon so thank you for the help. :). Also I've come to the conclusion that it's not a function. Would I be right?

OpenStudy (anonymous):

Let me look

OpenStudy (anonymous):

To determine if it is inverse you would take the Function f(x)= (x-8)^2 and use the Inverse function as your value for f(x) or x. If it equals x it is an inverse function. So we would get \[(\sqrt{x} + 8 - 8)^{2}\]

OpenStudy (anonymous):

Subtract 8 from 8 and you left with \[\sqrt{x}^{2}\] Which equals x. So \[y=\sqrt{x} + 8 ....is.. an.. inverse.. function of.. y = (x+8)^{2}\]

OpenStudy (anonymous):

Thank you.